# Fast Policy Learning for Linear-Quadratic Control with Entropy Regularization

### v3

@@ Abstract
This paper proposes and analyzes two new policy learning methods: regularized policy gradient (RPG) and iterative policy optimization (IPO), for a class of discounted linear-quadratic control (LQC) problems over an infinite time horizon with entropy regularization. Assuming access to the exact policy evaluation, both proposed approaches are proved to converge linearly in finding optimal policies of the regularized LQC. Moreover, the IPO method can achieve a super-linear convergence rate once it enters a local region around the optimal policy. Finally, when the optimal policy for an RL problem with a known environment is appropriately transferred as the initial policy to an RL problem with an unknown environment, the IPO method is shown to enable a super-linear convergence rate if the two environments are sufficiently close. Performances of these proposed algorithms are supported by numerical examples.
@@ 1 Introduction
Related works of policy gradient methods in LQC.
@@
Related works of entropy regularization.
@@
For system with infinite number of states and actions, the closest to our work in terms of model setup is [39]. Our work replaces their aggregated control setup with controls that are randomly sampled from the policy, which are more realistic in handling real-world systems. The focuses of these two papers are also different: theirs explained the exploitation–exploration trade-off with entropy regularization from a continuous-time stochastic control perspective and provided theoretical support for Gaussian exploration for LQC; while ours is on algorithms design and the convergence analysis for discrete-time dynamics. It is worth noting that recently [13] showed the linear convergence rate of the policy gradient method under the framework of [39] with aggregated controls in continuous time. To the best of our knowledge, our work is the first non-asymptotic convergence result for discrete-time LQC with randomized controls under entropy regularization. In addition, we establish the first local super-linear convergence result under the LQC setting.
Related works of transfer learning.
@@
Notations and organization.
Throughout the paper, we will denote, for any matrix $Z\in\mathbb{R}^{m\times d}$, $Z^{\top}$ for the transpose of $Z$, $\|Z\|$ for the spectral norm of $Z$, $\|Z\|_{F}$ for the Frobenius norm of $Z$, $\operatorname{Tr}(Z)$ for the trace of a square matrix $Z$, and $\sigma_{\min}(Z)$ (resp. $\sigma_{\max}(Z)$) for the minimum (resp. maximum) singular value of a square matrix $Z$. We will adopt $\mathcal{N}(\mu,\Sigma)$ for a Gaussian distribution with mean $\mu\in\mathbb{R}^{d}$ and covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$.
The rest of the paper is organized as follows. Section 2 introduces the problem and provides its theoretical solution using the dynamic programming principle. Section 3 presents the gradient dominance condition and related smoothness property. Section 4 introduces the RPG method and provides the global linear convergence result, and Section 5 proposes the IPO method, along with its global linear convergence and local super-linear convergence results. Section 6 shows that IPO leads to an efficient transfer learning scheme. Numerical examples are presented in Section 7.
Throughout this paper, proofs of main technical lemmas, unless otherwise specified, are deferred to Section 8.
@@ 2.1 Problem Formulation
Randomized policy and entropy regularization.
To enable entropy regularization for exploration in the context of learning, we focus on randomized Markovian policies that are stationary. Namely, define the admissible policy set as $\Pi:=\Big{\{}\pi:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{A})\Big{\}},$ with $\mathcal{X}$ the state space, $\mathcal{A}$ the action space, and $\mathcal{P}(\mathcal{A})$ the space of probability measures on action space $\mathcal{A}$. Here each admissible policy $\pi\in\Pi$ maps a state $x\in\mathcal{X}$ to a randomized action in $\mathcal{A}$.
For a given admissible policy $\pi\in\Pi$, the corresponding Shannon’s entropy is defined as
$\mathcal{H}(\pi(\cdot|x)):=-\int_{\mathcal{A}}\log\pi(u|x)d\pi(u|x),\qquad\forall x\in\mathcal{X}.$
@@
Objective function and dynamics.
@@
$J_{\pi}(x):=\mathbb{E}_{\pi}\left.\left[\sum_{t=0}^{\infty}\gamma^{t}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)\right|x_{0}=x\right],$ (2)
and for $t=0,1,2,\cdots$,
$x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\,x_{0}\sim\mathcal{D}.$ (3)
Here $x_{t}\in\mathcal{X}:=\mathbb{R}^{n}$ is the state of the system and the initial state $x_{0}$ follows an initial distribution $\mathcal{D}$. Here $u_{t}\in\mathcal{A}:=\mathbb{R}^{k}$ is the control at time $t$ following a policy $\pi$. In addition, $\{w_{t}\}_{t=0}^{\infty}$ are zero-mean independent and identically distributed (i.i.d) noises. We assume that $\{w_{t}\}_{t=0}^{\infty}$ have finite second moments. That is, $\operatorname{Tr}(W)<\infty$ with $W:=\mathbb{E}[w_{t}w_{t}^{\top}]$ for any $t=0,1,2,\cdots$. The matrices $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times k}$ define the system’s (transition) dynamics. $Q\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{k\times k}$ are matrices that parameterize the quadratic costs. $\gamma\in(0,1)$ denotes the discount factor and $\tau>0$ denotes the regularization parameter. The expectation in (2) is taken with respect to the control $u_{t}\sim\pi(\cdot|x_{t})$ and system noise $w_{t}$ for $t\geq 0$.
@@ 2.2 Optimal Value Function and Policy
While the optimal solution to the LQC problem is a well-explored topic, it is worth noting that, to the best of our knowledge, no prior work has presented a solution to the entropy-regularized LQC problem in the form of (2). Additionally, in the study of [39] for entropy-regularized LQC with continuous-time state dynamics, they focused on the state transitions with aggregated controls. This differs from the state transitions considered in (3), where the controls in the state transitions are randomly sampled from the policy $\pi$.
Optimal value function.
@@
The optimal value function $J^{*}:\mathcal{X}\rightarrow\mathbb{R}$ in (4) can be expressed as $J^{*}(x)=x^{\top}Px+q$ with $P$ and $q$ satisfying
$P$ $=Q+K^{*\top}RK^{*}+\gamma(A-BK^{*})^{\top}P(A-BK^{*}),$ (5)
$q$ $=\frac{1}{1-\gamma}\Big{[}\operatorname{Tr}(\Sigma^{*}(R+\gamma B^{\top}PB))-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma^{*}\right)\right)+\gamma\operatorname{Tr}(WP)\Big{]},$
where
$K^{*}=\gamma(R+\gamma B^{\top}PB)^{-1}B^{\top}PA,\quad\Sigma^{*}=\frac{\tau}{2}(R+\gamma B^{\top}PB)^{-1}.$ (6)
For any $x\in\mathcal{X}$, the corresponding optimal policy $\pi^{*}$ to (4) is a Gaussian policy: $\pi^{*}(\cdot|x)=\mathcal{N}(-K^{*}x,\Sigma^{*}).$
Proof of Theorem 2.1 relies on the following lemma, which establishes the optimal solution for the one-step reward function, given the presence of entropy regularization in the reward.
@@
For any given symmetric positive definite matrix $M\in\mathbb{R}^{k\times k}$ and vector $b\in\mathbb{R}^{k}$, the optimal solution $p^{*}\in\mathcal{P}(\mathcal{A})$ to the following optimization problem is a multivariate Gaussian distribution with covariance $\frac{\tau}{2}M^{-1}$ and mean $-\frac{1}{2}M^{-1}b$:
$\min_{p\in\mathcal{P}(\mathcal{A})}$ $\mathbb{E}_{u\sim p(\cdot)}\left[u^{\top}Mu+b^{\top}u+\tau\log p(u)\right],$
[…]
$p(u)\geq 0,\quad\forall u\in\mathcal{A}.$
Proof.
(of Theorem 2.1). By definition of $J^{*}$ in (4),
$J^{*}(x)=\min_{\pi\in\Pi}\mathbb{E}_{\pi}\Bigl{\{}x^{\top}Qx+u^{\top}Ru+\tau\log(\pi(u|x))+\gamma J^{*}(Ax+Bu+w)\Bigr{\}},$ (7)
where the expectation is taken with respect to $u\sim\pi(\cdot|x)$ and the noise term $w$, with mean 00 and covariance $W$. Stipulating
$J^{*}(x)=x^{\top}Px+q$ (8)
for a positive definite $P\in\mathbb{R}^{n\times n}$ and $q\in\mathbb{R}$ and plugging into (7), we can obtain the optimal value function with dynamic programming principle [4, 5, 15, 33]:
$J^{*}(x)$ $=x^{\top}Qx$
$\quad+\min_{\pi}\mathbb{E}_{\pi}\Bigl{\{}u^{\top}Ru+\tau\log(\pi(u|x))+\gamma\left[(Ax+Bu+w)^{\top}P(Ax+Bu+w)+q\right]\Bigr{\}}$
@@
$\quad+\min_{\pi}\mathbb{E}_{\pi}\Bigl{\{}u^{\top}(R+\gamma B^{\top}PB)u+\tau\log(\pi(u|x))+2\gamma u^{\top}B^{\top}PAx\Bigr{\}}.$ (9)
Now apply Lemma 2.2 to (9) with $M=R+\gamma B^{\top}PB$ and $b=2\gamma B^{\top}PAx$, one can get the optimal policy at state $x$:
$\pi^{*}(\cdot|x)$ $=\mathcal{N}\left(-\gamma(R+\gamma B^{\top}PB)^{-1}B^{\top}PAx,\frac{\tau}{2}(R+\gamma B^{\top}PB)^{-1}\right)$ (10)
$=\mathcal{N}\left(-K^{*}x,\Sigma^{*}\right),$ (11)
where $K^{*},\Sigma^{*}$ are defined in (6). To derive the associated optimal value function, we first calculate the negative entropy of policy $\pi^{*}$ at any state $x\in\mathcal{X}$:
[…]
Plug (10) and (12) into (9) to get
@@
$+\text{Tr}(\Sigma^{*}R)-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma^{*}\right)\right)+\gamma\left(\text{Tr}(\Sigma^{*}B^{\top}PB)+\text{Tr}(WP)+q\right).$
Combining this with (8), the proof is complete. ∎
@@ 3 Analysis of Value Function and Policy Gradient
Throughout the analysis, we assume that there exists $\rho\in(0,\frac{1}{\sqrt{\gamma}})$ satisfying $\|A-BK^{*}\|\leq\rho$ where $K^{*}$ is the optimal solution in Theorem 2.1. We consider the following domain $\Omega$ (i.e., the admissible control set) for $(K,\Sigma)$: $\Omega=\{K\in\mathbb{R}^{k\times n},\Sigma\in\mathbb{R}^{k\times k}:\|A-BK\|\leq\rho,\Sigma\succ 0,\Sigma^{\top}=\Sigma\}.$ For any $x_{0}\in\mathcal{X}$ following the initial distribution $\mathcal{D}$, we assume that $\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}x_{0}^{\top}]$ exists and $\mu:=\sigma_{\min}\left(\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}x_{0}^{\top}]\right)>0.$ For any $(K,\Sigma)\in\Omega$, define $S_{K,\Sigma}$ as the discounted state correlation matrix, i.e.,
$S_{K,\Sigma}:=\mathbb{E}_{\pi_{K,\Sigma}}\left[\sum_{i=0}^{\infty}\gamma^{i}x_{i}x_{i}^{\top}\right].$ (13)
@@
According to Theorem 2.1, the optimal policy of (1) is a Gaussian policy with a mean following a linear function of the state and a constant covariance matrix. In the remainder of the paper, we look for a parameterized policy of the form
$\pi_{K,\Sigma}(\cdot|x):=\mathcal{N}(-Kx,\Sigma),$ (14)
@@
for any $x\in\mathcal{X}$. With a slight abuse of notation, we use $J_{K,\Sigma}$ to denote $J_{\pi_{K,\Sigma}}$ and denote the objective in (1) as a function of $(K,\Sigma)$, given by
$C(K,\Sigma):=\mathbb{E}_{x\sim\mathcal{D}}\left[J_{K,\Sigma}(x)\right].$ (15)
@@
To analyze the dependence of $\Sigma$ in the objective function (15) for any fixed $K$, we also define $f_{K}:\mathbb{R}^{k\times k}\rightarrow\mathbb{R}$ as
@@
$f_{K}({\Sigma})=\frac{\tau}{2(1-\gamma)}\log\operatorname{det}({\Sigma})-\frac{1}{1-\gamma}\operatorname{Tr}\left({\Sigma}(R+\gamma B^{\top}P_{K}B)\right),\quad\forall\,{\Sigma}\succ 0.$ (16)
@@
By applying the Bellman equation, we can get $J_{K,\Sigma}(x)=x^{\top}P_{K}x+q_{K,\Sigma}$ with some symmetric positive definite matrix $P_{K}\in\mathbb{R}^{n\times n}$ and $q_{K,\Sigma}\in\mathbb{R}$ satisfying
$P_{K}$ $=Q+K^{\top}RK+\gamma(A-BK)^{\top}P_{K}(A-BK),$ (17)
$q_{K,\Sigma}$ $=\frac{1}{1-\gamma}\Big{[}\operatorname{Tr}(\Sigma(R+\gamma B^{\top}P_{K}B)-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma\right)\right)+\gamma\operatorname{Tr}(WP_{K})\Big{]}.$
@@
Note that (17) differs from equations (5) in terms of their solutions. In (5), the values of $K^{*}$ and $\Sigma^{*}$ are explicitly defined by $K^{*}=\gamma(R+\gamma B^{\top}PB)^{-1}B^{\top}PA$ and $\Sigma^{*}=\frac{\tau}{2}(R+\gamma B^{\top}PB)^{-1}$. By substituting these values into (5), one can obtain the solutions for $P$ and $q$, which define the optimal value function $J^{*}(x)=x^{\top}Px+q$.
@@
Meanwhile, $K$ and $\Sigma$ in (17) can take any admissible policy parameter values in $\Omega$, and the resulting $P_{K}$ and $q_{K,\Sigma}$ are functions of these policy parameters. The value function $J_{K,\Sigma}(x)$ derived from (17) represents the value starting from the state $x$ with the policy parameters $(K,\Sigma)$, which may or may not correspond to an optimal policy.
@@
We now provide an explicit form for the gradient of the cost function $C(K,\Sigma)$ with respect to $K$ and $\Sigma$. This explicit form will be used to show the gradient dominance condition in Lemma 3.2 and also in analyzing the algorithms in Sections 4 and 5.
@@
Take a policy in the form of (14) with parameter $(K,\Sigma)\in\Omega$, then the policy gradient has the following explicit form:
$\nabla_{K}C(K,\Sigma)$ $=2E_{K}S_{K,\Sigma},\quad\nabla_{\Sigma}C(K,\Sigma)=(1-\gamma)^{-1}\left(R-\frac{\tau}{2}\Sigma^{-1}+\gamma B^{\top}P_{K}B\right),$ (18)
where $E_{K}:=-\gamma B^{\top}P_{K}(A-BK)+RK$ and $S_{K,\Sigma}$ is defined in (13).
Gradient dominance.
@@
$C(K,\Sigma)-C(K^{*},\Sigma^{*})\leq\frac{\|S_{K^{*},\Sigma^{*}}\|}{4\mu^{2}\sigma_{\min}(R)}\|\nabla_{K}C(K,\Sigma)\|_{F}^{2}+\frac{1-\gamma}{\sigma_{\min}(R)}\|\nabla_{\Sigma}C(K,\Sigma)\|_{F}^{2}.$ (19)

### v4

@@ Abstract
This paper proposes and analyzes two new policy learning methods: regularized policy gradient (RPG) and iterative policy optimization (IPO), for a class of discounted linear-quadratic control (LQC) problems over an infinite time horizon with entropy regularization. Assuming access to the exact policy evaluation, both proposed approaches are proved to converge linearly in finding optimal policies of the regularized LQC. Moreover, the IPO method can achieve a super-linear convergence rate once it enters a local region around the optimal policy. Finally, when the optimal policy for an RL problem with a known environment is appropriately transferred as the initial policy to an RL problem with an unknown environment, the IPO method is shown to converge at a super-linear rate if the two environments are sufficiently close. A model-free version of the policy-based methods is also discussed. Performances of these proposed algorithms are supported by numerical examples.
keywords:
linear-quadratic control, reinforcement learning, policy gradient method, iterative policy optimization, entropy regularization, transfer learning
@@ 1 Introduction
Related works of policy gradient methods in LQC
@@
Related works of entropy regularization
@@
For system with infinite number of states and actions, the closest to our work in terms of model setup is [42]. Our work replaces their aggregated control setup with controls that are randomly sampled from the policy, which are more realistic in handling real-world systems. The focuses of these two papers are also different: theirs explained the exploitation–exploration trade-off with entropy regularization from a continuous-time stochastic control perspective and provided theoretical support for Gaussian exploration for LQC; while ours is on algorithms design and the convergence analysis. To the best of our knowledge, our work is the first non-asymptotic convergence result for LQC under entropy regularization.
Related works of transfer learning
@@
Notations and organization
Throughout the paper, we will denote, for any matrix $\in\mathbb{R}^{m\times d}, ZZ^{\top}$ for the transpose of $Z$, $\|Z\|$ for the spectral norm of $Z$, $\|Z\|_{F}$ for the Frobenius norm of $Z$, $\operatorname{Tr}(Z)$ for the trace of a square matrix $Z$, and $\sigma_{\min}(Z)$ (resp. $\sigma_{\max}(Z)$) for the minimum (resp. maximum) singular value of a square matrix $Z$. Let $\mathbb{S}^{d}_{+}$ denote the set of symmetric positive semi-definite matrices in $\mathbb{R}^{d\times d}$ and $\mathbb{S}^{d}_{++}$ for the subset of $\mathbb{S}^{d}_{+}$ consisting of symmetric positive definite matrices. We will adopt $\mathcal{N}(\mu,\Sigma)$ for a Gaussian distribution with mean $\mu\in\mathbb{R}^{d}$ and covariance matrix $\Sigma\in\mathbb{S}^{d}_{+}$.
The rest of the paper is organized as follows. Section 2 introduces the problem and provides its theoretical solution using the dynamic programming principle. Section 3 presents the gradient dominance condition and related smoothness property. Section 4 introduces the RPG method and provides the global linear convergence result, and Section 5 proposes the IPO method, along with its global linear convergence and local super-linear convergence results. Section 6 shows that IPO leads to an efficient transfer learning scheme. A model-free version of the policy-based method is discussed in Section 7. Numerical examples are presented in Section 8. Throughout this paper, proofs of main technical lemmas, unless otherwise specified, are deferred to Section 9.
@@ 2.1 Problem Formulation
Randomized policy and entropy regularization
To enable entropy regularization for exploration in the context of learning, we focus on randomized Markovian policies that are stationary. Namely, define the admissible policy set as $\Pi:=\{\pi:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{A})\}$, with $\mathcal{X}$ the state space, $\mathcal{A}$ the action space, and $\mathcal{P}(\mathcal{A})$ the space of probability measures on action space $\mathcal{A}$. Here each admissible policy $\pi\in\Pi$ maps a state $x\in\mathcal{X}$ to a randomized action in $\mathcal{A}$.
For a given admissible policy $\pi\in\Pi$, the corresponding Shannon’s entropy is defined as [19, 27]:
$\mathcal{H}(\pi(\cdot\mid x)):=-\int_{\mathcal{A}}\log\pi(u\mid x)\pi(u\mid x)\mathrm{d}u.$
@@
Objective function and dynamics
@@
(2.2) $J_{\pi}(x):=\mathbb{E}_{\pi}\left.\left[\sum_{t=0}^{\infty}\gamma^{t}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)\right|x_{0}=x\right],$
and such that for $t=0,1,2,\cdots$,
(2.3) $x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\,x_{0}\sim\mathcal{D}.$
Here $x_{t}\in\mathcal{X}:=\mathbb{R}^{n}$ is the state of the system and the initial state $x_{0}$ follows an initial distribution $\mathcal{D}$. Here $u_{t}\in\mathcal{A}:=\mathbb{R}^{k}$ is the control at time $t$ following a policy $\pi$. In addition, $\{w_{t}\}_{t=0}^{\infty}$ are zero-mean independent and identically distributed (i.i.d) noises. We assume that $\{w_{t}\}_{t=0}^{\infty}$ have finite second moments. That is, $\operatorname{Tr}(W)<\infty$ with $W:=\mathbb{E}[w_{t}w_{t}^{\top}]$ for any $t=0,1,2,\cdots$. The matrices $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times k}$ define the system’s (transition) dynamics. $Q\in{\color[rgb]{0,0,0}\mathbb{S}^{n}_{+}}$ and $R\in{\color[rgb]{0,0,0}\mathbb{S}_{++}^{k}}$ are matrices that parameterize the quadratic costs. $\gamma\in(0,1)$ denotes the discount factor and $\tau>0$ denotes the regularization parameter. The expectation in (2.1) is taken with respect to the control $u_{t}\sim\pi(\cdot|x_{t})$ and system noise $w_{t}$ for $t\geq 0$.
@@ 2.2 Optimal Value Function and Policy
While the optimal solution to the LQC problem is a well-explored topic, it is worth noting that, to the best of our knowledge, no prior work has presented a solution to the entropy-regularized LQC problem in the form of (2.1). Additionally, in the study of [42] for entropy-regularized LQC with continuous-time state dynamics, they focused on the state transitions with aggregated controls. This differs from the state transitions considered in (2.3), where the controls in the state transitions are randomly sampled from the policy $\pi$.
Optimal value function
@@
The optimal policy $\pi^{*}$ to (2.4) is a Gaussian policy: $\pi^{*}(\cdot|x)=\mathcal{N}(-K^{*}x,\Sigma^{*}),\forall x\in\mathcal{X},$ where
(2.5) $K^{*}=\gamma(R+\gamma B^{\top}PB)^{-1}B^{\top}PA,\quad\Sigma^{*}=\frac{\tau}{2}(R+\gamma B^{\top}PB)^{-1},$
with $P$ and $q$ satisfying
(2.6a) $P$ $=Q+K^{*\top}RK^{*}+\gamma(A-BK^{*})^{\top}P(A-BK^{*}),$
(2.6b) $q$ $=\frac{1}{1-\gamma}\Big(\operatorname{Tr}(\Sigma^{*}(R+\gamma B^{\top}PB))-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma^{*}\right)\right)+\gamma\operatorname{Tr}(WP)\Big).$
The optimal value function $J^{*}$ in (2.4) can be expressed as $J^{*}(x)=x^{\top}Px+q$.
Proof of Theorem 2.1 relies on the following lemma, which establishes the optimal solution for the one-step reward function with entropy regularization in the reward.
@@
For any given matrix $M\in\mathbb{S}^{k}_{+}$ and vector $b\in\mathbb{R}^{k}$, the optimal solution $p^{*}\in\mathcal{P}(\mathcal{A})$ to the following optimization problem (P) is a multivariate Gaussian distribution with covariance $\frac{\tau}{2}M^{-1}$ and mean $\frac{1}{2}M^{-1}b$:
(P) $\min_{\color[rgb]{0,0,0}p:\mathcal{A}\mapsto[0,\infty)}$ $\mathbb{E}_{u\sim p(\cdot)}\left[u^{\top}Mu+b^{\top}u+\tau\log p(u)\right],$
subject to $\int_{\mathcal{A}}p(u)du=1.$
Proof 2.3.
(of Theorem 2.1). By the definition of $J^{*}$ in (2.4),
(2.7) $J^{*}(x)=\min_{\pi\in\Pi}\mathbb{E}_{\pi}\Bigl\{x^{\top}Qx+u^{\top}Ru+\tau\log(\pi(u|x))+\gamma J^{*}(Ax+Bu+w)\Bigr\},$
where the expectation is taken with respect to $u\sim\pi(\cdot|x)$ and the noise term $w$, with mean 0 and covariance $W$. Stipulating
(2.8) $J^{*}(x)=x^{\top}Px+q$
for $P\in\mathbb{S}^{n}_{+}$ and $q\in\mathbb{R}$ and plugging into (2.7), we can obtain the optimal value function with dynamic programming principle [5, 7, 17, 37]:
$J^{*}(x)$ $=x^{\top}Qx+\min_{\pi}\mathbb{E}_{\pi}\Bigl\{u^{\top}Ru+\tau\log(\pi(u|x))$
$\quad+\gamma\left[(Ax+Bu+w)^{\top}P(Ax+Bu+w)+q\right]\Bigr\}$
@@
(2.9) $\quad+\min_{\pi}\mathbb{E}_{\pi}\Bigl\{u^{\top}(R+\gamma B^{\top}PB)u+\tau\log(\pi(u|x))+2\gamma u^{\top}B^{\top}PAx\Bigr\}.$
Now apply Lemma 2.2 to (2.9) with $M=R+\gamma B^{\top}PB$ and $b=2\gamma B^{\top}PAx$, one can get the optimal policy at state $x$:
(2.10) $\pi^{*}(\cdot|x)$ $=\mathcal{N}\left(-\gamma(R+\gamma B^{\top}PB)^{-1}B^{\top}PAx,\frac{\tau}{2}(R+\gamma B^{\top}PB)^{-1}\right)=\mathcal{N}\left(-K^{*}x,\Sigma^{*}\right),$
where $K^{*},\Sigma^{*}$ are defined in (2.5).
To derive the associated optimal value function, we first calculate the negative entropy of policy $\pi^{*}$ at any state $x\in\mathcal{X}$:
(2.11) $\mathbb{E}_{\pi^{*}}[\log(\pi^{*}(u|x))]=\int_{\mathcal{A}}\log(\pi^{*}(u|x))\pi^{*}(u|x)du=-\frac{1}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma^{*}\right)\right).$
Plug (2.10) and (2.11) into (2.9) to get
@@
$\quad+\text{Tr}(\Sigma^{*}R)-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma^{*}\right)\right)+\gamma\left(\text{Tr}(\Sigma^{*}B^{\top}PB)+\text{Tr}(WP)+q\right).$
Combining this with (2.8), we obtain the Riccati equation in (2.6a), which according to [6, Proposition 4.4.1] has a unique solution $P$. This is because (2.6a) takes the same form as the one for the classical LQR problem (without entropy regularization). With the unique solution $P$, we can then define the unique $q$ as in (2.6b), which finishes the proof.
@@ 3 Analysis of Value Function and Policy Gradient
Throughout the analysis, we assume that there exists $\rho\in(0,\frac{1}{\sqrt{\gamma}})$ satisfying $\|A-BK^{*}\|\leq\rho$ where $K^{*}$ is the optimal solution in Theorem 2.1. We consider the following domain $\Omega$ (i.e., the admissible control set) for $(K,\Sigma)$: $\Omega=\{K\in\mathbb{R}^{k\times n},\Sigma\in\mathbb{S}^{k}_{+}\}.$ For any $x_{0}\in\mathcal{X}$ following the initial distribution $\mathcal{D}$, we assume that $\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}x_{0}^{\top}]$ exists and $\mu:=\sigma_{\min}\left(\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}x_{0}^{\top}]\right)>0.$ For any $(K,\Sigma)\in\Omega$, define $S_{K,\Sigma}$ as the discounted state correlation matrix, i.e.,
(3.1) $S_{K,\Sigma}:=\mathbb{E}_{\pi_{K,\Sigma}}\left[\sum_{i=0}^{\infty}\gamma^{i}x_{i}x_{i}^{\top}\right].$
@@
According to Theorem 2.1, the optimal policy of (2.1) is a Gaussian policy with a mean following a linear function of the state and a constant covariance matrix. In the remainder of the paper, we look for a parameterized policy of the form
(3.2) $\pi_{K,\Sigma}(\cdot|x):=\mathcal{N}(-Kx,\Sigma),$
@@
for any $x\in\mathcal{X}$. With a slight abuse of notation, we use $J_{K,\Sigma}$ to denote $J_{\pi_{K,\Sigma}}$ and denote the objective in (2.1) as a function of $(K,\Sigma)$, given by
(3.3) $C(K,\Sigma):=\mathbb{E}_{x\sim\mathcal{D}}\left[J_{K,\Sigma}(x)\right].$
@@
To analyze the dependence of $\Sigma$ in the objective function (3.3) for any fixed $K$, we also define $f_{K}:\mathbb{R}^{k\times k}\rightarrow\mathbb{R}$ as
@@
(3.4) $f_{K}({\Sigma})=\frac{\tau}{2(1-\gamma)}\log\operatorname{det}({\Sigma})-\frac{1}{1-\gamma}\operatorname{Tr}\left({\Sigma}(R+\gamma B^{\top}P_{K}B)\right),\quad\forall\,{\Sigma}\succ 0.$
@@
By applying the Bellman equation, we can get $J_{K,\Sigma}(x)=x^{\top}P_{K}x+q_{K,\Sigma}$ with matrix $P_{K}\in\mathbb{S}^{n}_{+}$ and $q_{K,\Sigma}\in\mathbb{R}$ satisfying
(3.5) $P_{K}$ $=Q+K^{\top}RK+\gamma(A-BK)^{\top}P_{K}(A-BK),$
$q_{K,\Sigma}$ $=\frac{1}{1-\gamma}\Big(\operatorname{Tr}(\Sigma(R+\gamma B^{\top}P_{K}B)-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma\right)\right)+\gamma\operatorname{Tr}(WP_{K})\Big).$
@@
Note that (3.5) differs from equations (2.6a) in terms of their solutions. In (2.6a), the values of $K^{*}$ and $\Sigma^{*}$ are explicitly defined by $K^{*}=\gamma(R+\gamma B^{\top}PB)^{-1}B^{\top}PA$ and $\Sigma^{*}=\frac{\tau}{2}(R+\gamma B^{\top}PB)^{-1}$. By substituting these values into (2.6a), one can obtain the solutions for $P$ and $q$, which define the optimal value function $J^{*}(x)=x^{\top}Px+q$.
@@
Meanwhile, $K$ and $\Sigma$ in (3.5) can take any admissible policy parameter values in $\Omega$, and the resulting $P_{K}$ and $q_{K,\Sigma}$ are functions of these policy parameters. The value function $J_{K,\Sigma}(x)$ derived from (3.5) represents the value starting from state $x$ with policy parameters $(K,\Sigma)$, which may or may not correspond to an optimal policy.
@@
We now provide an explicit form for the gradient of the objective function $C(K,\Sigma)$ with respect to $K$ and $\Sigma$. This explicit form will be used to show the gradient dominance condition in Lemma 3.2 and also in analyzing the algorithms in Sections 4 and 5.
@@
Take a policy in the form of (3.2) with parameter $(K,\Sigma)\in\Omega$, then the policy gradient has an explicit form:
(3.6) $\nabla_{K}C(K,\Sigma)$ $=2E_{K}S_{K,\Sigma},\,\nabla_{\Sigma}C(K,\Sigma)=\frac{1}{1-\gamma}\left(R-\frac{\tau}{2}\Sigma^{-1}+\gamma B^{\top}P_{K}B\right),$
where $E_{K}:=-\gamma B^{\top}P_{K}(A-BK)+RK$ and $S_{K,\Sigma}$ is defined in (3.1).
Gradient dominance
@@
(3.7) $C(K,\Sigma)-C(K^{*},\Sigma^{*})\leq\frac{\|S_{K^{*},\Sigma^{*}}\|}{4\mu^{2}\sigma_{\min}(R)}\|\nabla_{K}C(K,\Sigma)\|_{F}^{2}+\frac{1-\gamma}{\sigma_{\min}(R)}\|\nabla_{\Sigma}C(K,\Sigma)\|_{F}^{2}.$